# Normalized solutions to mixed dispersion nonlinear Schrödinger system with coupled nonlinearity

### v1

@@ Abstract
where $\Delta^{2}u=\Delta(\Delta u)$ is the biharmonic operator, $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$, $N\geq 1$. $\rho^{2}$ stands for the prescribed mass, and $\lambda\in\mathbb{R}$ arises as a Lagrange multiplier. Such single constraint permits mass transformation in two materials. When $r_{1}+r_{2}\in\left(2,2+\frac{8}{N}\right]$, we obtain a dichotomy result for the existence of nontrivial ground states. Especially when $\alpha_{1}=\alpha_{2}$, the ground state exists for all $\rho>0$ if and only if $r_{1}+r_{2}<\min\left\{\max\left\{4,2+\frac{8}{N+1}\right\},2+\frac{8}{N}\right\}$. When $r_{1}+r_{2}\in\left(2+\frac{8}{N},\frac{2N}{(N-4)^{+}}\right)$ and $N\geq 2$, we obtain the existence of radial nontrivial mountain pass solution for small $\rho>0$.
Keywords: Biharmonic system, Normalized solutions, Ground state, Mountain pass solution.
@@ 1 Introduction
where $\alpha_{1}$, $\alpha_{2}$, $\beta$, $\rho>0$, $N\geq 1$, $r_{1}$, $r_{2}>1$, $r:=r_{1}+r_{2}\in(2,2^{**})$, $\Delta^{2}u=\Delta(\Delta u)$ is the biharmonic operator, and $\lambda\in\mathbb{R}$ arises as a Lagrange multiplier, which is unknown. Here
@@
is called Sobolev critical exponent. Any solution $(u,v,\lambda)$ of (1.1) satisfying (1.2) is usually called normalized solution.
In recent years, many researchers have considered the normalized solutions for the following single biharmonic equation
$\begin{cases}\Delta^{2}u+\alpha\Delta u+\lambda u=f(u)&\text{ in }\mathbb{R}^{N},\\ \int_{\mathbb{R}^{N}}u^{2}\mathrm{d}x=\rho^{2},\end{cases}$ (1.3)
where $\alpha\in\mathbb{R}$, $\rho>0$ and $\lambda$ is a Lagrange multiplier. In order to regularize and stabilize the solutions to the Schrödinger equation
@@
On the other hand, to prevent blow-up in finite time, Fibich, Ilan and Papanicolaou [9] also added a small fourth-order dispersion term in nonlinear Schrödinger equation as a nonparaxial correction. For more introduction about the background of (1.4), we refer to [1, 2].
@@
If $f(\psi)=e^{i\lambda t}f(u)$, up to a scaling, (1.4)-(1.5) with prescribed $L^{2}$-norm can be reduced to the equation (1.3). Note that the biharmonic term and Laplacian term are two dispersive terms, and a lack of homogeneity occurs provided $\alpha\neq 0$, which brings main difficulties in deriving the existence of solutions. One of approaches to obtain solutions of (1.3) is variational method. That is, the critical points of the associated energy functional
@@
$H^{2}(\mathbb{R}^{N}):=\left\{u\in L^{2}(\mathbb{R}^{N}):\nabla u,\Delta u\in L^{2}(\mathbb{R}^{N})\right\}$
endowed with the equivalent norm $\|u\|:=\left(\|\Delta u\|^{2}_{2}+\|u\|^{2}_{2}\right)^{\frac{1}{2}}$. Observe that the biharmonic term possesses dominating role in energy functional when $\alpha\leq 0$, while the Laplacian term has great effect on the geometry of energy functional when $\alpha>0$, especially which shows a convex-concave shape when we rescale the function $u$. Regardless of $\alpha\leq 0$ or $\alpha>0$, there exist a lot of works concentrated on the nonlinearities $f(u)=\mu|u|^{q-2}u+|u|^{p-2}u$ with $\mu\geq 0$.
The simplest case is the power nonlinearities $f(u)=|u|^{p-2}u$. Here we are convenient to introduce the parameter $\bar{r}:=2+\frac{8}{N}$, which is usually called $L^{2}$-critical exponent or mass critical exponent for fourth order equation. By Gagliardo-Nirenberg inequality (see Lemma 2.1 below) or dilations, the sign of $p-\bar{r}$ decides the geometry of the energy functional $E$ on $\Lambda_{\rho}$. Many researchers treated this type of problem according to the value of $p$ and the sign of $\alpha$.
@@
    $\overline{m}(\rho):=\underset{u\in\Lambda_{\rho}}{\inf}E(u)$
    is obviously sub-additive since the energy is invariant under translation. Whenever $\overline{m}(\rho)$ is negative, the vanishing of the minimizing sequence will not occur, hence $\overline{m}(\rho)$ can be attained. They proved that $\overline{m}(\rho)<0$ if $2<p<2+\frac{4}{N}$, so $\overline{m}(\rho)$ is attainable for all $\rho>0$. For $2+\frac{4}{N}\leq p<\bar{r}$, $\overline{m}(\rho)$ could be zero if the mass is large. As a result, the existence is a dichotomy result with respect to mass.
  • Bonheure et al. [2] investigated the case $\alpha<0$ and $p\in[\bar{r},2^{**})$. They proved the existence of the ground states when $\rho\in(c_{1},c_{2})$ for two numbers $c_{1}\geq 0,\,c_{2}\in(0,\infty]$ depending only on $N,\,p$. In particular, there holds that $c_{1}=0$ if $p>\bar{r}$. Moreover, they constructed minimax levels by $\mathbb{Z}_{2}$-genus, obtaining the multiplicity of radial normalized solutions.
  • When $\alpha>0$, the problem is more involved since the Laplacian has much effect on the shape of energy functional. Fernández et al. [8] established some non-homogeneous Gagliardo-Nirenberg inequalities, and showed some overall results about the existence and non-existence of global minimizers for $p\in(2,\bar{r}]$ and local minimizers for $p\in(\bar{r},2^{**})$. Furthermore, when $p\in(\bar{r},2^{**})$, Luo and Yang [14] proved that (1.3) has another mountain pass type solution as $\rho$ is small.
@@
As far as we are aware, it seems that there is no paper considering normalized solutions involving mixed dispersion nonlinear Schrödinger system. Motivated by above works, we are interested in the existence of normalized solutions for problem (1.1)-(1.2). That is, we search for $(u,v,\lambda)$ satisfying (1.1)-(1.2). A solution $(u,v,\lambda)$ is said nontrivial, which means that $u\neq 0$ and $v\neq 0$. The functional
@@
$S_{\rho}:=\left\{(u,v)\in H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}):\|u\|^{2}_{2}+\|v\|^{2}_{2}=\rho^{2}\right\}.$ (1.6)
To the best of our knowledge, very few works concerned the single constraint for $u$ and $v$ in nonlinear Schrödinger system. However the constraint (1.6) could be often encountered in physical world. In fact, $(u,v)\in S_{\rho}$ represents that the total mass of system is conserved, but the mass of $u$ and $v$ may transform mutually, where the transformation may be aroused due to chemical reactions or the movement in physical space.
Note that the energy functional $I$ under the different cases $r\in(2,\bar{r})$, $r=\bar{r}$ and $r\in(\bar{r},2^{**})$ has different geometry on $S_{\rho}$. When either $r\in(2,\bar{r})$ or $r=\bar{r}$ and $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$ ($\mathcal{D}_{1}$ is given in Theorem 1.1 below), we know that $I$ is coercive on $S_{\rho}$, see Lemma 3.1. Hence we concern whether the minimization problem
$m(\rho):=\inf_{(u,v)\in S_{\rho}}I(u,v)$ (1.7)
is attainable, whose minimizers are usually called ground states. As the method of [8, 15], the key ingredient is to establish the compactness of minimizing sequence. We emphasize here that $\alpha_{1}$, $\alpha_{2}>0$ and the fact that we can not obtain the strong convergence $\nabla u_{n}\rightarrow\nabla u$ in $L^{2}(\mathbb{R}^{N})$ only from $u_{n}\rightharpoonup u$ in $H^{2}(\mathbb{R}^{N})$ bring the main difficulties for problem (1.1)-(1.2). For that, we first establish a strict sub-additivity for $m(\rho)$ in Lemma 3.3. With this result and Lions’ concentration compactness principle, we provide in Theorem 3.5 an alternative result for minimizing sequence, that is, the minimizing sequence either vanishes or has a convergent subsequence.
@@
$J(u,v)=\frac{1}{2}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)-\frac{1}{2}\alpha_{1}\|\nabla u\|^{2}_{2}-\frac{1}{2}\alpha_{2}\|\nabla v\|^{2}_{2},\quad\forall(u,v)\in S_{\rho}.$
It can be seen that $J$ is coercive on $S_{\rho}$, so $m^{J}(\rho)$ is always well-defined. We will show in Lemma 3.2 that $m^{J}(\rho)$ is never achieved. If the minimizing sequence for $m(\rho)$ vanishes, there must hold $m(\rho)=m^{J}(\rho)$. Therefore we can rule out the vanishing of the minimizing sequence if $m(\rho)<m^{J}(\rho)$. So the key step is to establish the comparison between $m(\rho)$ and $m^{J}(\rho)$. As the ideas in [8], this comparison can be reduced into whether the supremum of
@@
Assume that $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$. Let $\bar{r}=2+\frac{8}{N}$, $r=r_{1}+r_{2}$ and $\mathcal{D}_{1}=\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}{C}_{N,r}^{r}$, where ${C}_{N,r}$ is the optimal constant for Gagliardo-Nirenberg inequality (2.1). Then there exists some
$\rho^{*}\in\begin{cases}[0,\infty)\quad&\mbox{if}\;\;r<\bar{r},\\ \left[0,\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}\right)\quad&\mbox{if}\;\;r=\bar{r},\end{cases}$
@@
$\rho\in\begin{cases}(\rho^{*},\infty)\quad&\mbox{if}\;\;r<\bar{r},\\ \left(\rho^{*},\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}\right)\quad&\mbox{if}\;\;r=\bar{r},\end{cases}$
@@
From the proof of Theorem 1.1, the strict subadditivity is important for obtaining the existence of solutions. While considering
$\begin{cases}\Delta^{2}u+\alpha_{1}\Delta u+\lambda_{1}u=\beta r_{1}|u|^{r_{1}-2}|v|^{r_{2}}u&\text{ in }\mathbb{R}^{N},\\ \Delta^{2}v+\alpha_{2}\Delta v+\lambda_{2}v=\beta r_{2}|u|^{r_{1}}|v|^{r_{2}-2}v&\text{ in }\mathbb{R}^{N},\end{cases}$
with two constraints
$\int_{\mathbb{R}^{N}}u^{2}\mathrm{d}x=\rho^{2}_{1},\leavevmode\nobreak\ \int_{\mathbb{R}^{N}}v^{2}\mathrm{d}x=\rho^{2}_{2}\quad\text{with}\leavevmode\nobreak\ \rho_{1},\leavevmode\nobreak\ \rho_{2}>0,$
the existence issue is very different from single constraint case, since it could not establish the strict subadditivity if we proceed as the proof of Lemma 3.3.
Notice that $\rho^{*}$ in Theorem 1.1 has chance to be 0, this means that $m(\rho)$ can be achieved for all $\rho>0$. Next, we aim to find the borderline to guarantee $\rho^{*}=0$. To this end, the cases that $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$ have great differences. Actually the problem in the case $\alpha_{1}=\alpha_{2}$ is very similar to the single equation
We will in subsection 3.2 make use of the estimates of [8] to give a borderline for $\rho^{*}=0$. However, when $\alpha_{1}\neq\alpha_{2}$, the estimates are challenging and elusive. We only know that the same condition as the case $\alpha_{1}=\alpha_{2}$ can guarantee $\rho^{*}>0$, but it is difficult for us to conclude how $\rho^{*}$ equals 0. Our result is the following.
Theorem 1.3.
Assume $\alpha_{1},\,\alpha_{2},\,\beta>0$ and $r\leq\bar{r}$. Let $\rho^{*}$ be given in Theorem 1.1. Then $\rho^{*}>0$ if
$\max\left\{4,2+\frac{8}{N+1}\right\}\leq r\leq 2+\frac{8}{N}.$ (1.10)
In particular, when $\alpha_{1}=\alpha_{2}$, (1.10) is also a necessary condition for $\rho^{*}>0$.
Remark 1.4.
@@
Theorems 1.1 and 1.3 can tell us that when $\alpha_{1}=\alpha_{2}$, the nontrivial ground state exists for all $\rho>0$ if and only if
$2<r_{1}+r_{2}<\min\left\{\max\left\{4,2+\frac{8}{N+1}\right\},2+\frac{8}{N}\right\}.$
@@
Now, we consider the case $r\in(\bar{r},2^{**})$. Observe that when the mass is small, the energy functional $I$ possesses a mountain pass geometry. More precisely, when $\beta\rho^{r-2}<c^{*}(N,r)$, there exist $R_{1}>R_{0}>0$ such that for any $c\in(R_{0},R_{1})$, $I$ has a positive lower bound on $S_{\rho}$ with the restrict $\|\Delta u\|_{2}^{2}+\|\Delta v\|_{2}^{2}=c^{2}$, where
$c^{*}(N,r):=\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{r\gamma_{r}-2}{(r\gamma_{r}-1)\max\{\alpha_{1},\alpha_{2}\}}\right)^{r\gamma_{r}-2},$
@@
$\gamma_{r}:=\frac{N(r-2)}{4r}$ and $\mathcal{D}_{1}=\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}{C}_{N,r}^{r}$. On the other hand, we can find two points $(u_{0},v_{0})$, $(u_{1},v_{1})\in S_{\rho}$ such that $I(u_{0},v_{0}),\,I(u_{1},v_{1})<0$ with $\|\Delta u_{0}\|_{2}^{2}+\|\Delta v_{0}\|_{2}^{2}<R_{0}^{2}$ and $\|\Delta u_{1}\|_{2}^{2}+\|\Delta v_{1}\|_{2}^{2}>R_{1}^{2}$. Thus, we can establish a mountain pass structure, precisely see Section 4. Using Jeanjean’s method in [10], we can obtain a Palais-Smale sequence $\{(u_{n},v_{n})\}$ approaching Pohožaev manifold, but whose compactness is very involved. By virtue of Lagrange multiplier rule, we derive a sequence of Lagrange multipliers $\{\lambda_{n}\}$ corresponding to the sequence $\{(u_{n},v_{n})\}$. For the compactness, a key step is to deduce $\underset{n\to\infty}{\liminf}\,\lambda_{n}>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$. Following the method in [14], if in addition $\beta\rho^{r-2}<c_{*}(N,r)$, we can reach this aim, where
@@
$c_{*}(N,r)$ :=$&\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{1-\gamma_{r}}{\gamma_{r}}\frac{4}{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}\right)^{\frac{r\gamma_{r}-2}{2}}\quad\quad\quad\text{ if }\frac{1}{2}<\gamma_{r}<1,\\ &\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{r-2}{2(r\gamma_{r}-1)}\frac{4}{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}\right)^{\frac{r\gamma_{r}-2}{2}}\leavevmode\nobreak\ \quad\text{ if }0<\gamma_{r}\leq\frac{1}{2}.\end{aligned}\right.$
@@
Let $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$, $N\geq 2$ and $r:=r_{1}+r_{2}\in(\bar{r},2^{**})$. Then for any $\rho>0$ satisfying

### v2

@@ Abstract
where $\Delta^{2}u=\Delta(\Delta u)$ is the biharmonic operator, $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$, $N\geq 1$. $\rho^{2}$ stands for the prescribed mass, and $\lambda\in\mathbb{R}$ arises as a Lagrange multiplier. Such single constraint permits mass transformation in two materials. When $r_{1}+r_{2}\leq 2+\frac{8}{N}$, we obtain a dichotomy result with respect to the mass for the existence of nontrivial ground states. Especially when $\alpha_{1}=\alpha_{2}$, the ground state exists for all $\rho>0$ if and only if $r_{1}+r_{2}<\min\left\{\max\left\{4,2+\frac{8}{N+1}\right\},2+\frac{8}{N}\right\}$. When $r_{1}+r_{2}\in\left(2+\frac{8}{N},\frac{2N}{(N-4)^{+}}\right)$ and $N\geq 2$, we obtain the existence of radial nontrivial mountain pass solution for sufficiently small $\rho>0$.
Keywords: Biharmonic operator, Nonlinear Schrödinger system, Normalized solutions, Ground state, Mountain pass solution.
@@ 1 Introduction
where $\alpha_{1}$, $\alpha_{2}$, $\beta$, $\rho>0$, $N\geq 1$, $r_{1}$, $_{2}>1, r:=r$$r_{1}+r_{2}<2^{**}$, $\Delta^{2}u=\Delta(\Delta u)$ is the biharmonic operator, and $\lambda\in\mathbb{R}$ arises as a Lagrange multiplier, which is unknown. Here
@@
which is called Sobolev critical exponent when $N\geq 5$. Any solution $(u,v,\lambda)$ of (1.1) satisfying (1.2) is usually called a normalized solution of (1.1).
In recent years, many researchers have considered the normalized solutions for the following biharmonic Schrödinger equation
[…]
where $\alpha\in\mathbb{R}$, $\rho>0$ and $\lambda$ is a Lagrange multiplier. There is a fruitful research background on the biharmonic nonlinear Schrödinger equation. On the one hand, in order to regularize and stabilize the solutions to the Schrödinger equation
@@
On the other hand, to prevent blow-up in finite time, Fibich, Ilan and Papanicolaou [9] added a small fourth-order dispersion term in nonlinear Schrödinger equation as a nonparaxial correction. We omit the details about these introductions, one can refer to [1, 2].
@@
If $f(\psi)=e^{i\lambda t}f(u)$, up to a scaling, (1.4)-(1.5) with prescribed $L^{2}$-norm can be reduced to the equation (1.3). Note that the biharmonic term and Laplacian term are two dispersive terms, and a lack of homogeneity occurs provided $\alpha\neq 0$, which brings the main difficulties in deriving the existence of solutions. An approach to obtain solutions of (1.3) is variational method. That is, the critical points of the associated energy functional
@@
$H^{2}(\mathbb{R}^{N})=\left\{u\in L^{2}(\mathbb{R}^{N}):\nabla u,\Delta u\in L^{2}(\mathbb{R}^{N})\right\}$
endowed with the equivalent norm $\|u\|:=\left(\|\Delta u\|^{2}_{2}+\|u\|^{2}_{2}\right)^{\frac{1}{2}}$. Observe that the biharmonic term $\|\Delta u\|_{2}^{2}$ possesses dominating role in energy functional when $\alpha\leq 0$, while the Laplacian term has great effect on the geometry of energy functional when $\alpha>0$, especially which shows a convex-concave shape when we rescale the function $u$. Regardless of $\alpha\leq 0$ or $\alpha>0$, there exist a lot of works concentrated on the nonlinearities $f(u)=\mu|u|^{q-2}u+|u|^{p-2}u$ with $\mu\geq 0$.
The simplest case is the power nonlinearities $f(u)=|u|^{p-2}u$. Consequently the energy functional is
$\widetilde{E}(u)=\frac{1}{2}\|\Delta u\|_{2}^{2}-\frac{\alpha}{2}\|\nabla u\|_{2}^{2}-\frac{1}{p}\int_{\mathbb{R}^{N}}|u|^{p}\mathrm{d}x,\quad\forall u\in\Lambda_{\rho}.$
It is convenient to introduce the parameter $\bar{r}:=2+\frac{8}{N}$, which is usually called $L^{2}$-critical exponent or mass critical exponent for fourth order equation. By Gagliardo-Nirenberg inequality (see Lemma 2.1 below) and the inequality (2.2), or by dilations, the sign of $p-\bar{r}$ decides the geometry of the energy functional $\widetilde{E}|_{\Lambda_{\rho}}$. Observe that $\widetilde{E}$ is coercive on $\Lambda_{\rho}$ if $2<p<\bar{r}$; and $\widetilde{E}$ is unbounded from below on $\Lambda_{\rho}$ if $\bar{r}<p\leq 2^{**}$.
Many researchers treated the problem (1.3) with $f(u)=|u|^{p-2}u$ according to the value of $p$ and the sign of $\alpha$.
@@
    $\overline{m}(\rho):=\underset{u\in\Lambda_{\rho}}{\inf}\widetilde{E}(u)$
    is obviously sub-additive since the energy is invariant under translation. Whenever $\overline{m}(\rho)$ is negative, the vanishing of the minimizing sequence will not occur, hence $\overline{m}(\rho)$ can be attained. They proved that $\overline{m}(\rho)<0$ if $2<p<2+\frac{4}{N}$, so $\overline{m}(\rho)$ is attainable for all $\rho>0$. For $2+\frac{4}{N}\leq p<\bar{r}$, $\overline{m}(\rho)$ could be zero if the mass is large. As a result, they obtained a dichotomy result with respect to mass.
  • Bonheure et al. [2] investigated the case $\alpha<0$ and $p\in[\bar{r},2^{**})$. They proved the existence of the ground states when $\rho\in(c_{1},c_{2})$ for two numbers $c_{1}\geq 0,\,c_{2}\in(c_{1},\infty]$ depending only on N,pN,\,p. In particular, there holds that $c_{1}=0$ if $p>\bar{r}$. Moreover, they constructed minimax levels by $\mathbb{Z}_{2}$-genus, obtaining the multiplicity of radial normalized solutions.
  • When $\alpha>0$, the problem is more involved since the Laplacian has an effect on the shape of energy functional. Fernández et al. [8] established some non-homogeneous Gagliardo-Nirenberg inequalities, and showed some results about the existence and non-existence of global minimizers for $p\in(2,\bar{r}]$ and local minimizers for $p\in(\bar{r},2^{**})$. Furthermore, when $p\in(\bar{r},2^{**})$, Luo and Yang [14] proved that (1.3) has another mountain pass type solution as $\rho$ is small.
@@
As far as we are aware, it seems that there is no literature considering normalized solutions involving mixed dispersion nonlinear Schrödinger system. Motivated by the above works, we are interested in the existence of normalized solutions for problem (1.1)-(1.2). That is, we search for $(u,v,\lambda)$ satisfying (1.1)-(1.2). A solution $(u,v,\lambda)$ is said nontrivial, which means that $u\neq 0$ and $v\neq 0$. The functional
@@
$S_{\rho}=\left\{(u,v)\in H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N}):\|u\|^{2}_{2}+\|v\|^{2}_{2}=\rho^{2}\right\}.$ (1.6)
To the best of our knowledge, very few works concerned the single constraint for $u$ and $v$ in nonlinear Schrödinger system. However the constraint (1.6) could be often encountered in physical world. In fact, $(u,v)\in S_{\rho}$ represents that the total mass of a system is conserved, but the mass of $u$ and $v$ may transform mutually, where the transformation may be aroused due to chemical reactions or the movement in physical space.
In this paper, we are focused on $\alpha_{1},\alpha_{2},\beta>0$. Note that the energy functional $I$ under the different cases $\in(2,\bar{r}), r=r$$\bar{r}$ and $r\in(\bar{r},2^{**})$ has different geometry on $S_{\rho}$. If either $r\in(2,\bar{r})$ or $r=\bar{r}$ and $\rho<\left(\frac{1}{2\mathcal{D}_{1}\beta}\right)^{\frac{N}{8}}$ where
$\mathcal{D}_{1}=\left(\frac{r_{1}}{r}\right)^{r_{1}}\left(\frac{r_{2}}{r}\right)^{r_{2}}{C}_{N,r}^{r},$ (1.7)
and ${C}_{N,r}$ is the optimal constant for Gagliardo-Nirenberg inequality (2.1), we get in Lemma 3.1 that $I$ is coercive on $S_{\rho}$. Hence we ask whether the minimization problem
$m(\rho)=\inf_{(u,v)\in S_{\rho}}I(u,v)$ (1.8)
is attainable. Usually, normalized solutions with least energy are called normalized ground states (for short, we omit the word “normalized”), so any minimizer of (1.8) is a ground state. As the method of [8, 15], the key ingredient is to establish the compactness of minimizing sequence. We emphasize here that $\alpha_{1}$, $\alpha_{2}>0$ and the fact that one can not obtain the strong convergence $\nabla u_{n}\rightarrow\nabla u$ in $L^{2}(\mathbb{R}^{N})$ only from $u_{n}\rightharpoonup u$ in $H^{2}(\mathbb{R}^{N})$ brings the main difficulties for problem (1.1)-(1.2). For that, we first establish a strict sub-additivity for $m(\rho)$ in Lemma 3.3. With this result and Lions’ concentration compactness principle, we provide in Theorem 3.5 an alternative result for minimizing sequence, that is, the minimizing sequence either vanishes or has a convergent subsequence.
@@
$J(u,v)=\frac{1}{2}\left(\|\Delta u\|^{2}_{2}+\|\Delta v\|^{2}_{2}\right)-\frac{\alpha_{1}}{2}\|\nabla u\|^{2}_{2}-\frac{\alpha_{2}}{2}\|\nabla v\|^{2}_{2},\quad\forall(u,v)\in S_{\rho}.$
It can be seen that $J$ is coercive on $S_{\rho}$, so $m^{J}(\rho)$ is always well-defined. We will show in Lemma 3.2 that $m^{J}(\rho)$ is never achieved. If the minimizing sequence for $m(\rho)$ vanishes, there must hold $m(\rho)=m^{J}(\rho)$. Therefore we can rule out the occurrence of vanishing if $m(\rho)<m^{J}(\rho)$. So the key step is to establish the comparison between $m(\rho)$ and $m^{J}(\rho)$. As the ideas in [8], this comparison can be reduced into whether the supremum of
@@
Assume that $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$. Let $\bar{r}=2+\frac{8}{N}$, $r=r_{1}+r_{2}$ and $\mathcal{D}_{1}$ be given in (1.7). Then there exists some
[…]
@@
[…]
@@
Notice that it is possible that $\rho^{*}$ in Theorem 1.1 is equal to 0, which means that $m(\rho)$ can be achieved for all $\rho>0$. Next, we aim to find the borderline to guarantee $\rho^{*}=0$. To this end, the cases that $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}\neq\alpha_{2}$ have great differences. Actually the problem in the case $\alpha_{1}=\alpha_{2}$ is very similar to the single equation
We will in subsection 3.2 make use of the estimates of [8] to give a borderline for $\rho^{*}=0$. However, when $\alpha_{1}\neq\alpha_{2}$, the estimates are challenging and elusive. In our paper, for $\alpha_{1}\neq\alpha_{2}$, we obtain the condition that can guarantee $\rho^{*}>0$, which is similar as the case $\alpha_{1}=\alpha_{2}$, but it is difficult for us to conclude how $\rho^{*}$ equals 0. Our result is the following.
Theorem 1.2.
Assume $\alpha_{1},\,\alpha_{2},\,\beta>0$ and $r\leq 2+\frac{8}{N}$. Let $\rho^{*}$ be given in Theorem 1.1. Then $\rho^{*}>0$ if
$\max\left\{4,2+\frac{8}{N+1}\right\}\leq r\leq 2+\frac{8}{N}.$ (1.11)
Furthermore, when $\alpha_{1}=\alpha_{2}$, (1.11) is also a necessary condition for $\rho^{*}>0$.
@@
Precisely, under the assumptions that $r_{1},r_{2}>1$ and $r_{1}+r_{2}\leq 2+\frac{8}{N}$, Theorems 1.1 and 1.2 say that when $\alpha_{1}=\alpha_{2}$, the nontrivial ground state exists for all $\rho>0$ if and only if
$r_{1}+r_{2}<\min\left\{\max\left\{4,2+\frac{8}{N+1}\right\},2+\frac{8}{N}\right\}.$
Now, we consider the case $r\in(\bar{r},2^{**})$. For the convenience of compactness, we will work in a subspace of $H^{2}(\mathbb{R}^{N})\times H^{2}(\mathbb{R}^{N})$, which consists of all radial functions, it is denoted by $H^{2}_{r}(\mathbb{R}^{N})\times H^{2}_{r}(\mathbb{R}^{N})$ where
$H^{2}_{r}(\mathbb{R}^{N})=\left\{u\in H^{2}(\mathbb{R}^{N}):u\mbox{ is radially symmetric}\right\}.$ (1.12)
@@
Observe that when the mass is small, the energy functional $I$ possesses a mountain pass structure. More precisely, when $\beta\rho^{r-2}<c^{*}(N,r)$, there exist $R_{1}>R_{0}>0$ such that for any $c\in(R_{0},R_{1})$, $I$ has a positive lower bound on $S_{\rho}$ with the restriction $\|\Delta u\|_{2}^{2}+\|\Delta v\|_{2}^{2}=c^{2}$, where
$c^{*}(N,r)=\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{r\gamma_{r}-2}{(r\gamma_{r}-1)\max\{\alpha_{1},\alpha_{2}\}}\right)^{r\gamma_{r}-2},$
@@
$\gamma_{r}:=\frac{N(r-2)}{4r}$ and $\mathcal{D}_{1}$ is given in (1.7). On the other hand, we can find two points $(u_{0},v_{0})$, $(u_{1},v_{1})\in S_{\rho}$ such that $I(u_{0},v_{0}),\,I(u_{1},v_{1})<0$ with respectively $\|\Delta u_{0}\|_{2}^{2}+\|\Delta v_{0}\|_{2}^{2}<R_{0}^{2}$ and $\|\Delta u_{1}\|_{2}^{2}+\|\Delta v_{1}\|_{2}^{2}>R_{1}^{2}$. Thus, we can establish a mountain pass structure, precisely see Section 4. Using Jeanjean’s method in [10], one can obtain a Palais-Smale sequence $\{(u_{n},v_{n})\}$ approaching Pohožaev manifold, namely
$2\int_{\mathbb{R}^{N}}(|\Delta u_{n}|^{2}+|\Delta v_{n}|^{2})\mathrm{d}x-\alpha_{1}\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}\mathrm{d}x-\alpha_{2}\int_{\mathbb{R}^{N}}|\nabla v_{n}|^{2}\mathrm{d}x-2\beta r\gamma_{r}\int_{\mathbb{R}^{N}}|u_{n}|^{r_{1}}|v_{n}|^{r_{2}}\mathrm{d}x=o(1),$ (1.13)
and by virtue of Lagrange multiplier rule, we derive a sequence of Lagrange multipliers $\{\lambda_{n}\}$ corresponding to the Palais-Smale sequence $\{(u_{n},v_{n})\}$. That is
As the usual process, it is not difficult to show that the sequence $(u_{n},v_{n})$ is bounded in $H_{r}^{2}(\mathbb{R}^{N})\times H_{r}^{2}(\mathbb{R}^{N})$, but whose compactness is very involved. As previously emphasized, the main difficulty is that we can not obtain the strong convergence $\nabla u_{n}\rightarrow\nabla u$ in $L^{2}(\mathbb{R}^{N})$ only from the weak convergence $u_{n}\rightharpoonup u$ in $H_{r}^{2}(\mathbb{R}^{N})$. Motivated by Luo and Yang [14], we multiply (1.14) by $(u_{n},v_{n})$, and combining (1.13), we can cancel the common terms $\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}\mathrm{d}x$ and $\int_{\mathbb{R}^{N}}|\nabla v_{n}|^{2}\mathrm{d}x$. Thus the compactness could escape from dealing with the convergence of $\nabla u_{n}$ and $\nabla v_{n}$ in $L^{2}$. Moreover, if $(u_{n},v_{n})$ is not compact in $H^{2}_{r}(\mathbb{R}^{N})\times H^{2}_{r}(\mathbb{R}^{N})$, up to a subsequence, we can deduce $\underset{n\to\infty}{\lim}\,\lambda_{n}\leq\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}$, see (4.12). Naturally, a key step to get the compactness is to obtain
$\underset{n\to\infty}{\lim}\,\lambda_{n}>\frac{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}{4}.$ (1.15)
In Section 4, we will prove that (1.15) can holds if $\beta\rho^{r-2}<c_{*}(N,r)$ where
@@
$c_{*}(N,r)$ $:=\left\{\begin{aligned} &\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{1-\gamma_{r}}{\gamma_{r}}\frac{4}{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}\right)^{\frac{r\gamma_{r}-2}{2}}\quad\quad\quad\text{ if }\frac{1}{2}<\gamma_{r}<1,\\ &\frac{1}{2\mathcal{D}_{1}(r\gamma_{r}-1)}\left(\frac{r-2}{2(r\gamma_{r}-1)}\frac{4}{\max\{\alpha^{2}_{1},\alpha^{2}_{2}\}}\right)^{\frac{r\gamma_{r}-2}{2}}\penalty 10000\ \quad\text{ if }0<\gamma_{r}\leq\frac{1}{2}.\end{aligned}\right$.
@@
Let $\alpha_{1}$, $\alpha_{2}$, $\beta>0$, $r_{1}$, $r_{2}>1$, $N\geq 2$ and $r\in(\bar{r},2^{**})$. Then for any $\rho>0$ satisfying